7. For circular motion of radius 15 cm and 400 g mass the time period is 0.15 s. What is its centripetal acceleration?
Answer #1

Given,

The radius of the circular path, R\ =\ 15\ cm

=\ 0.15\ m

Mass, M\ =\ 400\ g

=\ 0.4\ kg

Time taken to complete one rotation, t\ =\ 0.15\ s

The angle subtended in one rotation, \theta\ =\ 2\pi

Angular velocity, \omega\ =\ \dfrac{\theta}{t}

  =\ \dfrac{2\pi}{0.15}\ rad/s

=\ 41.87\ rad/s

Centripetal acceleration,a\ =\ \omega^2 R

=\ (41.87)^2\times 0.15

=\ 262.92\ m/s^2